OpenStudy (anonymous):

Find the equation in standard form which passes through (–2, –7) and (1, –1). You must show all work. i got -2x+7=-5 can someone help me

OpenStudy (anonymous):

how'd you come upon this answer?

OpenStudy (chaise):

You form should be written in: y=mx+c There is no y in this equation :s?

OpenStudy (anonymous):

i use the point slope formula

OpenStudy (anonymous):

point slope formula states \[y-y_1=m(x-x_1)\] you'll always have a y and x with a coefficient

OpenStudy (anonymous):

yah so is what i got correct

OpenStudy (anonymous):

no because your y in the equation never came out

OpenStudy (anonymous):

what did you get for your slope

OpenStudy (anonymous):

i got 2

OpenStudy (anonymous):

correct now using one of those points with m=2 you get \[y-(-1)=2(x-1)

OpenStudy (anonymous):

\[y-(-1)=2(x-1)\]

OpenStudy (anonymous):

\[y+1=2x-2\]

OpenStudy (anonymous):

should i plug in my point (1,-1) into the y=mx +b formula to get y=2x-3 which in standard form is y-2x=-3

OpenStudy (anonymous):

mx+b is slope intercept form you can only use this if you know the y intercept

OpenStudy (anonymous):

you need to use the point slope formula

OpenStudy (anonymous):

oh okay i accidentally messed up the 2(x-1) i missed the negative and that messed me up. So the answer would be y-2x=-3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much!!